OpenStudy (anonymous):

Cars A and B are travelling 30 and 40 miles per hour, respectively. If car B starts out 5 miles behind car A, how many hours will it take for car B to overtake car A?

OpenStudy (anonymous):

Solve the following equation for t: 30 t + 5 = 40 t t=1/2 hour

OpenStudy (anonymous):

Now for this problem the distance from the origin be x at which b overtake a then x x - 5 -- = -------- 40 30 which will give x = 20 now calculate time which will be 20/40 = 1/2hr